) A financial department has three printers, A, B and C, which print at different speeds. Documents are routed to the first available printer. The probabilities that a document is routed to printers A, B or C are 0.6, 0.3 and 0.1, respectively. Occasionally a printer will jam and destroy a printout. The probability that printers A, B and C will jam are 0.01, 0.05 and 0.04, respectively. (a) Find the probability that a printout is destroyed when a printer jams.

2) Given the print out is destroyed by a printer jam, What is the probability that printer A is involved
Answer #1

L +9) gwen P(A) =0.6 - (B) 0.3, P(C)=0. PJ|A) FO. O P(JB) 0.05, PC510) = 0.04 S: Jamed & pointout is destroyed P(+) 7 PCI/A).